OpenStudy (anonymous):

Please help! Will give a medal and become a fan of best answer! Please show all steps. :) What is the quotient? State any restrictions on the variable. 9-x^2/2x^2+5x-3 divided by 5x-15/5x-5

OpenStudy (anonymous):

((((9−( x2 2 )(x2))+5x)−( 3 5 )(x))−( 15 5 )(x))−5 =9+(−0.5x4)+5x+(−0.6x)+(−3x)+(−5) =(−0.5x4)+1.4x+4 =−0.5x4+1.4x+4

OpenStudy (anonymous):

Please fan and like if this helped. :D

OpenStudy (anonymous):

The steps will help but it looks all broken up. Let me see if I can figure out what goes where

OpenStudy (anonymous):

oK

OpenStudy (anonymous):

((((9−(x22)(x2))+5x)−(35)(x))−(155)(x))−5=9+(−0.5x4)+5x+(−0.6x)+(−3x)+(−5)=(−0.5x4)+1.4x+4=−0.5x4+1.4x+4

jimthompson5910 (jim_thompson5910):

When you divide fractions, you flip the second fraction and multiply. From there, you factor as much as possible and reduce. This is all shown in the attached PDF.